OpenStudy (anonymous):

Find the inverse of the function f(x) = (x - 4)/(x + 4) x does not equal -4

OpenStudy (anonymous):

let y=(x-4)/(x+4) 4y +xy=x-4. 4(y+1)=x(1-y). x=4(y+1)/(1-y). Therefore,the inverse of the given function is F(x)=4(x+1)/(1-x) Where x not equal to 1/